solving xy''+2y'+xy=0 by power series
1
Expert's answer
2019-05-08T15:17:11-0400
"y(x)=\sum_{n=0}^\infin a_nx^n"

"y'(x)=\sum_{n=1}^\infin na_nx^{n-1}"

"y''(x)=\sum_{n=2}^\infin n(n-1)a_nx^{n-2}"

"x\sum_{n=2}^\infin n(n-1)a_nx^{n-2}+2\sum_{n=1}^\infin na_nx^{n-1}+x\sum_{n=0}^\infin a_nx^n=0"

"\sum_{n=2}^\infin n(n-1)a_nx^{n-1}+2\sum_{n=1}^\infin na_nx^{n-1}+\sum_{n=0}^\infin a_nx^{n+1}=0"

"\sum_{n=2}^\infin n(n-1)a_nx^{n-1}+2\sum_{n=1}^\infin na_nx^{n-1}+\sum_{n=2}^\infin a_{n-2}x^{n-1}=0"

"2a_1+\sum_{n=2}^\infin [n(n-1)a_n+2na_n+a_{n-2}]x^{n-1}=0"

"a_1=0"

For

"n\ge2"

"n(n-1)a_n+2na_n+a_{n-2}=0"

"a_n=-a_{n-2}\/[(n(n-1)+2n]=-a_{n-2}\/[n(n-1)]"

So


"a_2=-a_0\/6=-a_0\/(2*3)"

"a_3=a_5=a_7=a_9=...=0"

"a_4=-a_2\/20=a_0\/120=a_0\/(2*3*4*5)"

"a_6=-a_4\/42=-a_0\/(2*3*4*5*6*7)"

"y(x)=a_0(1-x^2\/(2*3)+x^4\/(2*3*4*5)-x^6\/(2*3*4*5*6*7)+(-1)^kx^{2k}\/(2k+1)!)"

"y(x)=a_0 \sum_{k=0}^ \infin(-1)^kx^{2k}/(2k+1)!"